You are here: Foswiki>AdaptiveOptics Web>Logs>Plan_201202 (23 Feb 2012, JuanGuerra)Edit Attach

AO/Pisces Technical Observing Plan

Prior to Feb 3rd (ie Daytime):

DONE Test placement of ND holder. (Doug O., JC)

Michael can install the holder ASAP. Michael should place a piece of aluminum tape on one quadrant of the ND holder so the 90 degree corner is at the center of the ND holder.

=> Waiting for confirmation from Don that we can proceed.

From last email we can proceed to power on the computer and set the web cam and the lamp (+cardboard box)

The holder target is not in place. Talk with Doug in the mountain if he can take out the baffle and place the target and put it back.

DONE Closed dome test of ND holder position (JC)

Once Pisces is cooled, with closed dome we can take images with Pisces in the J band and hopefully see the dark aluminum tape quadrant. Then Michael (or Instrument crew) can move ND holder to put the 90 degree corner in the center of Pisces image. This will require Doug or JC to take Pisces images and give instructions for movement (interact with the mountain staff (Doug O). Be sure to record direction of ND holder motions relative to motion on Pisces images. (Michael and JC)

This job requires to activate Pisces, lights on the dome and telescope at zenith. Also Pisces has to be cold.

  • Not Conclusive Strange image quality on Pisces makes it difficult to see anything

ALERT! On-sky vignetting of ND holder position (JC)

=> Find an appropriate crowded field to check for vignetting on edges of Pisces image. (Open Cluster ) M36 or M37 can ne used at the beggining of the night.

=> JC prepare to perform photometry. Quick photometry.

* Not Possible with Strange Pisces images

DONE Pisces Web Camera (Doug O., JC)

Mount web camera so we can see Filter Wheel encoder

=> Waiting for confirmation from Don that we can proceed. From last Don email (1st on February) we can proceed to the installation of the webcam and the lamp.

DONE Cool Pisces (Doug O.)

Feb 1st. (Doug started to pump up and cool down) (completed 3PM)

DONE Pisces Computer (Doug O.)

  • Power up computer and CCD controller (Michael, now Doug), (contact person in the mountain from the 1st is Doug ...)
  • Test communication to Pisces (Doug)

=> Waiting for cooling of Pisces and confirmation from Don that we can proceed

Ok from the last Don email .

DONE Upgrade Observe routines (Doug)

  1. Install new preset telescope modes (IT#3738)
    • ADAPTIVETTM_TRACK, ADAPTIVETTM_GUIDE, ADAPTIVETTM_ACTIVE
    • ADAPTIVEACE_TRACK, ADAPTIVEACE_GUIDE, ADAPTIVEACE_ACTIVE
    • ADAPTIVEICE_TRACK, ADAPTIVEICE_GUIDE, ADAPTIVEICE_ACTIVE
  2. Install Position Angle Calculator *Currently we run a perl script written by Steve Allanson. I have a routine to calculate this same position angle so that we do not spill LN2. I need to install it into the Observe routines, install some logic (choose PA for a cardinal directions) and ask user, in realtime, if this value is fine, or allow input
  3. Install new RunAO commands (IT#3537)
    • NORMAL, REPOINT, SKIPREF, SKIPSTART, SKIPSTARTREPOINT, SKIPALL
  4. Make sure that at the end of a dither pattern a 0,0 offset is sent

DONE Create needed catalogs (JC)

Science catalogues from;

  • We have catalogs from John.
  • We have coordinates from Ian.
  • Communicating with Chick.
  • Get ADI test object from Laird.

ALERT! Prepare Observing scripts (JC, Doug)

Prepare technical observing scripts, incorporating late February targets if possible.

DONE Test Power W Unit (JC, Doug)

Simone and friends are going to startup AGw#2 and tests the pyramid sensor and all parts of the W before Thursday. Doug and JC will rerun startup with Simone and friends when they arrive on the mountain on Thursday or Friday.

ALERT! Use AO/Pisces exposure time calculator (JC)

ALERT! Test AO QuickLook routines (JC, Doug)

ALERT! Feb AO observing proposals (JC, Doug)

Get late Feb AO observing proposals so we can taylor our AO/Pisces night to test the need observing moves

  • Get details from John
    • Standard dither patterns
    • Not ADI
    • Several filters (JHK) but no ND's
    • AO References stars from 14 to 17.6 mag

  • Chick
    • ADI and Standard dither patterns
    • FeII filter
    • ND3 + J, H, Ks
    • AO Reference star ?? mag

  • Test ADI mode for Laird
    • ADI mode
    • ND3 + J, H, Ks
    • Find star to use for test (Laird)

  • Ian
    • High strehl not required (10-20)
    • J0946
      • AO Reference star 14.4 R mag
      • 18" from target
      • Target ~ 21 mag/arcsec^2 in K
      • ~ 3 hours open shutter in Ks
    • ULAS 1120+0641
      • AO Reference stars 16.2 R mag
      • 15" from target
      • Target K ~ 18.2 estimated
      • ~ 10 minutes open shutter in Ks

ALERT! ND filter usage

Laird: ND3 + J, H, Ks Chick: ND3 + J, H, Ks

Nighttime:

ALERT! Collect sky flats (JC)

Perform skyflat with and without ND during evening and morning twilight.

Observations occurs during the second half of the night, poor sky conditions (cloudy)

ALERT! First closed AO loop (All)

Close the loop with a simple case bright object (Rmag 8 or 9 ). Close the AO loop on bright star with ND filter so we are ready to determine focus offsets immediately after.

DONE Test new TCS flags (Doug)

  • New Preset flag
    • Send MODE=ADAPTIVEACE_TRACK preset with ROTMODE=PARALLACTIC PA=#
    • Send RunAO: should close the AO loop without doing guiding and active optics step
  • Test New RunAO flag
    • Send normal MODE=ADAPTIVEACE_ACTIVE preset
    • Stop AO loop on engineering GUI
    • IIF will report Preset Canceled
    • Start the AO loop from the engineering GUI
    • Send RunAO SkipAll to force IIF to make preset active again.

ALERT! Focus Offsets (All)

  • Determine W focus offset for J, H, Ks filters with ND3.
  • Arcetri person or people will be awake and help determine focus offsets.
  • Should do this early in the night so they can get to bed for their daytime work.

ALERT! Check for Vignetting by ND holder (JC)

  • Remove ND and check to see if there is any vignetting at the edges.
  • Look at crowded field.
  • JC will perform photometry to determine if dimming on the edges.

ALERT! Set Static Offset in PCS (Doug)

Set static offsets for pointing so target lands near center of Pisces image on preset. Currently center of rotation (target position) is near the left edge of Pisces.

  • Tested and works. Need to get proper values. Current values put on-axis star in upper left quadrant

ALERT! Science-like Observing (All)

Test science observations. Should test the situations which programs for late Feb partner observing. The goal is to accomplish:
  • Test AO closed loop on several types of object. (JC)
    • very bright - Close
    • very faint - Hill
  • ADI mode (Doug)
  • Test upgraded Observe routines. (Doug)
  • Use Quick-Look routines in realtime (All)

Other Tests
  • Check 13 Hz vibration by lowering M1 (both) while AO -loop is closed. (20 minutes)

-- DougMiller - 31 Jan 2012
Topic revision: r7 - 23 Feb 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback